# Topological Optimal Transport for Geometric Cycle Matching

### v1

@@ Abstract
Topological data analysis is a powerful tool for describing topological signatures in real world data. An important challenge in topological data analysis is matching significant topological signals across distinct systems. In geometry and probability theory, optimal transport formalises notions of distance and matchings between distributions and structured objects. We propose to combine these approaches, constructing a mathematical framework for optimal transport-based matchings of topological features. Building upon recent advances in the domains of persistent homology and optimal transport for hypergraphs, we develop a transport-based methodology for topological data processing. We define measure topological networks, which integrate both geometric and topological information about a system, introduce a distance on the space of these objects, and study its metric properties, showing that it induces a geodesic metric space of non-negative curvature. The resulting Topological Optimal Transport (TpOT) framework provides a transport model on point clouds that minimises topological distortion while simultaneously yielding a geometrically informed matching between persistent homology cycles.
@@ 1 Introduction
Topological data analysis (TDA) is a quickly growing field in computational and applied topology. In recent years, TDA has established itself as an effective framework to analyse, cluster, and detect patterns in complex data [52]. One of the key algorithms in TDA is persistent homology (PH), a computational tool that describes the structure of data based on topological features persisting across different scales [35, 18, 22]. In a nutshell, the PH algorithm proceeds by building a nested sequence of discrete spaces describing the input data at increasingly coarse scales, known as a filtration of simplicial complexes. The topological features of this filtration are then quantified by computing the homology groups of each simplicial complex in the sequence. A structural theorem [56] guarantees that the birth, death, and evolution of homology classes in the filtration can be summarised as a multi-set, called a persistence diagram (PD). PDs have been shown to contain rich information about the initial data, and PH-based data analysis approaches appear in an ever-increasing number of applications across multiple fields in modern science [40, 42, 51, 44].
@@
Optimal transport (OT) is a far-reaching mathematical theory that, in its most basic and classical formulation, formalises the problem of finding a matching of a given probability distribution to another as efficiently as possible, in terms of a cost function [50, 36]. When the cost function is induced by the distance function on a metric space $(X,d)$, optimal transport provides a natural lifting of the ground metric to a metric on the space of probability distributions supported on $X$, commonly known as the Wasserstein distance. An extension of the optimal transportation problem which has recently received considerable attention is the Gromov-Wasserstein (GW) problem [31, 32, 8, 43], in which the probability distributions to be matched are defined over different metric spaces, rather than supported on a common space. The objective is then to find matchings that minimise distortion of pairwise distances. While the GW distance was originally introduced to compare metric measure spaces and can be viewed as a relaxation of the well known Gromov-Hausdorff distance [31], some of its most successful applications address the problem of matching more general structured objects such as (labelled) graphs [54, 53, 10, 48]. Among its numerous advantages, the GW distance has brought substantial improvements to graph processing tasks over classical methods, and can be efficiently approximated by a range of numerical algorithms in practice [53, 37, 20, 9]. On the theoretical front, the space of gauged measure spaces (i.e., a measure space $(X,\mu)$ together with a symmetric gauge function $k:X\times X\to\mathbb{R}$) endowed with the GW distance has the geometry of an Alexandrov space, and a Riemannian orbifold structure can be developed in this setting [43, 10]. More recently, inspired by the goal of encoding relations in complex systems as higher-order networks, a variant of the GW problem, referred to as co-optimal transport [46], was applied to model hypergraphs both from a theoretical and applied point of view [11].
@@
Figure 1: The Topological Optimal Transport (TpOT) problem. From left to right: two input point clouds, their persistence diagrams, and corresponding PH-hypergraphs. The objective of the TpOT problem combines a Gromov-Wasserstein distortion on the point clouds (geometric information), (partial) Wasserstein matching of points in the persistence diagrams (topological information) and HyperCOT on the PH-hypergraphs (coupling of geometric and topological information).
@@
In this paper, we combine the TDA and OT approaches described above to develop a transport-based theory for topological data processing. Our approach is inspired by recent extensions built upon persistent homology computations: the hyperTDA framework [2] encodes each point in a persistence diagram, together with a geometric realisation of its homology class as a generating cycle, as a higher-order network called a PH-hypergraph. In a PH-hypergraph, vertices corresponds to data points used as input to the PH computation, and hyperedges are given by the generators of features in the persistence diagram. We define the topological optimal transport (TpOT) problem, by considering a trade-off between preservation of geometric relationships, preservation of topological features, and coupling the two via the PH-hypergraph structure (see Figure 1). The output is (1) a matching between points that is topologically driven and geometrically informed, coupled with (2) a matching between persistent homology classes that is geometrically driven and topologically informed.
@@
Finding meaningful ways of matching topological features across distinct systems is an important challenge in TDA. The past few years have seen considerable effort in addressing this question, with a number of different solutions proposed [13, 3, 25, 38, 55, 21], many of which bridge between topological data analysis and optimal transport [23, 27, 30]. A key aspect of the TpOT method is leveraging existing transportation theory for persistence diagrams to drive a matching between topological features, while also taking into account geometric information and the spatial interconnectivity of homology generators. By endowing PH-hypergraphs with probability measures, we introduce the concept of measure topological networks, and we formalise a framework for studying the TpOT induced distance $d_{\mathrm{TpOT},p}$. Our main contributions are as follows:
@@
  • We develop a flexible measure-theoretic formalism for simultaneously encoding the geometry and topology of a finite point cloud, in the form of the aforementioned measure topological networks, as well as a family of distances $d_{\mathrm{TpOT},p}$, $p\geq 1$, between these objects; see Definitions 1 and 2, respectively.
@@
  • We show that the metric induced by $d_{\mathrm{TpOT},p}$ is geodesic, characterise the exact form of the geodesics and use this characterisation to show that this metric space is non-negatively curved, in the sense of Alexandrov; see Theorems 2, 3 and 4, respectively.
  • Our framework is centred on an optimisation problem for which we provide efficient numerical algorithms, which we demonstrate on a variety of examples; see Section 5.
@@ 2 Mathematical background
Our formalisation of the TpOT problem relies on constructions in topological data analysis and measure theory. The aim of this section is to outline the mathematical background needed to define measure topological networks and the family of distances $d_{\mathrm{TpOT},p}$ described in the introduction.
@@ 2.1 Persistent homology and persistent homology-hypergraphs
Let $(X,d)$ be a finite metric space – for example, a point cloud $X=\{x_{1},\cdots,x_{N}\}\subset\mathbb{R}^{n}$ as illustrated in Figure 2A. For any $\varepsilon>0$, the Vietoris-Rips complex $K(X)_{\varepsilon}$ (see e.g. [18, Ch.III.2]) is the simplicial complex obtained from $X$ by adding a $k$-simplex $[x_{i_{0}},x_{i_{1}},\cdots,x_{i_{k}}]$ whenever the distance between all pairs of points in $\{x_{i_{0}},x_{i_{1}},\cdots,x_{i_{k}}\}$ is less than $\varepsilon$. Note that $K(X)_{\varepsilon}$ is a sub-complex of $K(X)_{\varepsilon^{\prime}}$ whenever $\varepsilon\leq\varepsilon^{\prime}$. Thus, as $\varepsilon$ grows, this yields a nested sequence of simplicial complexes:
$\mathcal{K}_{\rm VR}(X):=K(X)_{\varepsilon_{0}}\hookrightarrow K(X)_{\varepsilon_{1}}\hookrightarrow\cdots\hookrightarrow K(X)_{\varepsilon_{M}}.$
@@
Here, the numbers $\varepsilon_{i}$, $i=1,\ldots,M$, are the parameters corresponding to the creation of new simplices; see Figure 2B.
@@
Figure 2: Computing PH and the PH-hypergraph (A) A point cloud and (B) a filtration of simplicial complexes built on it. (C) The $1$-dimensional persistence diagram of the filtration in (B). (D) The PH-hypergraph constructed after computing a representative cycle (red cycles in (B)) for each class in (C).
@@
More generally, for a finite metric space $X$, a filtration over $X$ is a finite sequence $\mathcal{K}=(K_{\varepsilon_{1}},\ldots,K_{\varepsilon_{M}})$ of simplicial complexes, indexed by an increasing sequence of real numbers $\varepsilon_{i}\leq\varepsilon_{i+1}$, together with simplicial maps $K_{\varepsilon_{i}}\to K_{\varepsilon_{i+1}}$ for all $i$, such that the vertex set of each $K_{\varepsilon_{i}}$ is a subset of $X$. Common alternatives to the Vietoris-Rips complex in the literature with a similar interpretation include the Čech complex [18, Ch.III.2] and the Alpha complex [18, Ch.III.4]. Clearly, different constructions of the filtration $\mathcal{K}$ will influence the results of the persistent homology computation and their interpretation.
@@
A filtration $\mathcal{K}$ over $X$ can be analysed by computing the simplicial homology groups of each simplicial complex in the sequence and the corresponding induced maps. Collectively, this information is called the persistent homology (PH) of $\mathcal{K}$. The Structural Theorem [57], guarantees that, when using coefficients in $\mathbb{Z}/2\mathbb{Z}$, the birth, death, and evolution of homology classes in a filtration of simplicial complexes can be summarised in a multi-set, called a persistence diagram. We use $\Lambda$ to denote the upper-diagonal cone in the upper-half plane, $\Lambda=\{(x,y)\in\mathbb{R}^{2},y>x>0\}$. The persistence diagram $D$ associated to the filtration $\mathcal{K}$ is then a collection of points in $\Lambda$: $D=\{(b_{i},d_{i})\}_{i=1}^{|D|}\subset\Lambda$.
@@
Each point $(b_{i},d_{i})\in D$ represents a homology class in the filtration $\mathcal{K}$, with coordinates corresponding to the birth $b_{i}$ and death $d_{i}$ parameters of the class (see Figure 2C). The persistence $d_{i}-b_{i}>0$ of a homology class is a measure of significance of the underlying topological feature. The persistence diagram $D$ can be interpreted as a topological descriptor of $X$. A local, geometric realisation of each homology class can be obtained by computing a representative cycle generating the class: see for instance the red cycles in Figure 2B. Computing cycles depends on a number of choices [5, 16, 34, 49, 29], and therefore different representatives may lead to differing downstream results.
@@
The recently introduced PH-hypergraph [2] encodes the algebraic information of a persistence diagram, together with the local geometric interpretation of generating cycles, in a single higher-order network. Recall that a hypergraph is defined by a pair $(V,E)$ consisting of a set of vertices $V$ and a set of hyperedges $E$. Given a point cloud $X$, we may construct an associated filtration $\mathcal{K}$ together with a choice $g=\{c_{1},\ldots,c_{|D|}\}$ of generating cycles for classes in the corresponding persistence diagram $D$. These data can be summarised in a PH-hypergraph $H=H(X,\mathcal{K},g)$ in which we take $X$ to be the vertex set, and one hyperedge added for each cycle in $g$ (see Figure 2D). Moreover, for every cycle $c\in g$, a point $x_{i}\in X$ belongs to the corresponding hyperedge $e(c)$ if it is a vertex of any simplex in $c$. While the hypergraph structure $H$ clearly depends on the choice of generators $g$ and the choice of filtration $\mathcal{K}$, for a fixed filtration $\mathcal{K}$ the qualitative (hyper)network structure of the PH-hypergraph has been empirically shown to be stable to noise and different choices of generating cycles [2].
@@ 2.2 Optimal transport on persistence diagrams
Consider now two Borel probability measures $\mu$ and $\mu^{\prime}$ supported on a Polish metric space $(X,d)$. For $p\in[1,\infty)$, the $p$-Wasserstein distance [50, Definition 6.1] between $\mu$ and $\mu^{\prime}$ is then defined as
$d_{\mathrm{W},p}(\mu,\mu^{\prime}):=\left(\inf_{\pi\in\Pi(\mu,\mu^{\prime})}\int_{X\times X}d(x,y)^{p}\>\mathrm{d}\pi(x,y)\right)^{1/p}.$ (1)
This metric is the central object of study in the optimal transport (OT) literature. It is a standard result that the infimum in (1) is realised (i.e. it is in fact a minimum) and a minimiser $\pi$ of (1) is referred to as an optimal coupling. Intuitively, this is a probabilistic matching between points in the supports of $\mu$ and $\mu^{\prime}$.
The optimal transport approach of finding matchings has inspired different notions of distance on the space of persistence diagrams: the most popular construction is known as the bottleneck distance [12], which is related to the $\infty$-Wasserstein distance. In this article, we are more interested in similar adaptations of the $p$-Wasserstein distances for $1\leq p<\infty$. The usual modification of the Wasserstein matching problem to compare persistence diagrams is to allow points to be matched to a virtual point $\partial_{\Lambda}$, representing the diagonal line $\{(a,a)\mid a\in\mathbb{R}\}\subset\mathbb{R}^{2}$ (i.e. homology classes with vanishing persistence). Formally, for $p\in[1,\infty)$, the $p$-Wasserstein distance between two persistence diagrams $D$ and $D^{\prime}$ (see [28, Equation 6]) is defined as
$d_{\mathrm{W},p}^{\mathrm{PD}}(D,D^{\prime})^{p}=\min_{\pi\in\Pi(D,D^{\prime})}\left(\sum_{(a,b)\in\pi}\|a-b\|^{p}_{p}+\sum_{s\in U_{\pi}}\|s-\operatorname{Proj}_{\partial_{\Lambda}}(s)\|^{p}_{p}\right),$ (2)
@@
  • Slightly abusing notation, to illustrate the analogy to the OT Wasserstein distance, we denote by $\Pi(D,D^{\prime})$ the set of all partial matchings between points representing homology classes in $D$ and $D^{\prime}$. For any $\pi\in\Pi(D,D^{\prime})$, $\pi\subset D\times D^{\prime}$ has the property that each $a\in D$ and $b\in D^{\prime}$ appear in at most one ordered pair in $\pi$.
@@
The output from solving (2) is a partial matching $\pi$ between homology classes in $D$ and $D^{\prime}$ which is optimal with respect to the cost induced by the $\|\cdot\|_{p}$-norm in $\mathbb{R}^{2}$, restricted to $\Lambda$. By construction, this matching only depends on the coordinates of points in the persistence diagram that correspond to homology classes. Notably, the matching is agnostic to geometric and topological aspects of the data that are not captured in the persistence diagram representations. While the generality of the persistence diagram representation may contribute to its wide applicability, in many contexts this may be a limitation.
Observe that the notation and concepts used to define $d_{\mathrm{W},p}^{\mathrm{PD}}$ are similar to those that appear in the optimal transport framework described above; indeed, the family of persistence diagram distances (2) can be generalised via the language of partial OT [17]. We define a measure persistence diagram to be a Radon measure $\mu$ supported on $\Lambda$. Then, any finite persistence diagram $D=\{(b_{i},d_{i})\}_{i=1}^{|D|}$ determines a measure persistence diagram:

### v2

@@ Abstract
Topological data analysis is a powerful tool for describing topological signatures in real world data. An important challenge in topological data analysis is the task of matching significant topological signals across distinct systems. Optimal transportation provides a geometric and probabilistic approach to formalising notions of distances and matchings between measures and structured objects more generally. Building upon recent advances in the domains of persistent homology and optimal transport for hypergraphs, we develop an approach for finding geometrically and topologically informed matchings between point cloud datasets and their topological features. We define measure topological networks, which integrate both geometric and topological information about a system, and introduce a distance on the space of these objects. Our distance is defined in terms of a Topological Optimal Transport (TpOT) problem which seeks to transport mass that preserves geometric as well as topological relations, in the sense of minimising the induced geometric and topological distortions. We study the metric properties of this distance and show that it induces a geodesic metric space of non-negative curvature. We demonstrate our approach in a number of numerical experiments.
@@ 1 Introduction
Topological data analysis (TDA) is a quickly growing field in computational and applied topology. In recent years, TDA has established itself as an effective framework to analyse, cluster, and detect patterns in complex data [63]. One of the key algorithms in TDA is persistent homology (PH), a computational tool that describes the structure of data based on topological features persisting across different scales [45, 21, 26]. In brief, PH computation proceeds by building a nested sequence of discrete spaces describing the input data at increasingly coarse scales, known as a filtration of simplicial complexes. The topological features of this filtration are then quantified by computing the homology groups of each simplicial complex in the sequence. A structural theorem [67] guarantees that the birth, death, and evolution of homology classes in the filtration can be summarised as a multi-set, called a persistence diagram (PD). PDs have been shown to contain rich information about the initial data, and PH-based data analysis approaches appear in an ever-increasing number of applications across multiple fields in modern science [50, 53, 62, 55].
@@
Optimal transport (OT) is a mathematical framework that, in its classical formulation, formalises the problem of finding a matching between two given probability distributions that is optimal in terms of a prescribed cost [61, 46]. When the cost function is induced by the distance between points in a metric space $(X,d)$, optimal transport provides a natural lifting of the ground metric to a metric on the space of probability distributions supported on $X$, commonly known as the Wasserstein distance (see for instance, [61, Theorem 6.18]). An extension of the optimal transportation problem which has recently received considerable attention is the Gromov-Wasserstein (GW) problem [41, 42, 9, 54], in which the probability distributions to be matched are defined over different metric spaces, rather than supported on a common space. The objective is then to find matchings that minimise distortion of pairwise distances. While the GW distance was originally introduced to compare metric measure spaces and can be viewed as a relaxation of the well known Gromov-Hausdorff distance [41], it has seen recent success in matching more general structured objects such as (labelled) graphs [65, 64, 11, 58]. Among other advantages, GW distances and matchings have brought substantial improvements to graph processing tasks over classical methods, and can be efficiently approximated by a range of numerical algorithms in practice [64, 47, 24, 10]. On the theoretical front, the space of gauged measure spaces (i.e., measure spaces $(X,\mu)$ together with a symmetric gauge function $k:X\times X\to\mathbb{R}$) endowed with the GW distance has the geometry of an Alexandrov space, and a Riemannian orbifold structure can be developed in this setting [54, 11]. More recently, inspired by the goal of encoding relations in complex systems as higher-order networks, a variant of the GW problem, referred to as co-optimal transport [49], was applied to model hypergraphs both from a theoretical and applied point of view [13].
@@
Figure 1: Topological Optimal Transport (TpOT). From left to right: (i) two input point clouds, (ii) their persistence diagrams, and (iii) corresponding PH-hypergraphs. The objective of the TpOT problem shares information across these three levels via a combination of distortion functionals: a Gromov-Wasserstein distortion on the point clouds (geometric information), (partial) Wasserstein matching of points in the persistence diagrams (topological information) and HyperCOT on the PH-hypergraphs (coupling of geometric and topological information).
@@
In this paper, we combine the TDA and OT approaches described above to develop a transport-based method for topological data processing. Our approach is inspired by recent extensions built upon persistent homology computations: the hyperTDA framework [2] encodes each persistence feature – represented as a point in a persistence diagram – together with a geometric realisation of its corresponding homology class as a generating cycle in the input point cloud. The resulting object is a higher-order network called a PH-hypergraph. In a PH-hypergraph, vertices correspond to data points used as input to the PH computation, and hyperedges are given by the generators of features in the persistence diagram. We define the topological optimal transport (TpOT) problem, by considering a trade-off between preservation of geometric relationships (encoded in pairwise proximity of points in each point cloud), preservation of topological features (encoded in the PH-hypergraph), and a coupling of the two via the PH-hypergraph structure (see Figure 1). The output is (1) a matching between points that is geometrically driven and topologically informed, together with (2) a matching between persistent homology classes that is topologically driven and geometrically informed.
@@
Finding meaningful ways to match topological features across distinct systems is an important challenge in TDA. In recent years there has been considerable effort towards addressing this question, with a number of different solutions proposed [15, 3, 30, 48, 66, 25], many of which bridge the topological data analysis and optimal transport literatures [27, 34, 39]. A key aspect of the TpOT method is leveraging existing approaches for optimal matchings of persistence diagrams to drive a matching between topological features, while also taking into account geometric information and the spatial connectivity of homology generators. By endowing PH-hypergraphs with probability measures, we introduce measure topological networks as a structure for representing both geometric and PH information, and we formalise a framework for studying the TpOT induced distance $d_{\mathrm{TpOT},p}$.
Our contributions are as follows:
@@
  • We develop a flexible measure-theoretic formalism for simultaneously encoding the geometry and topology of a point cloud, in the form of measure topological networks, as well as a family of distances $d_{\mathrm{TpOT},p},p\geq 1$, between these objects; see Definitions 1 and 3, respectively.
@@
  • We show that the metric induced by $d_{\mathrm{TpOT},p}$ is geodesic, characterise the exact form of the geodesics, and use this characterisation to show that this metric space is non-negatively curved in the sense of Alexandrov; see Theorems 6, 7 and 8, respectively.
  • Our framework is formulated in terms of a non-convex optimisation problem, and we provide numerical algorithms for its approximate solution. We demonstrate its application in practice on a variety of examples; see Section 5.
@@ 2 Mathematical background
Our formalisation of the TpOT problem relies on constructions in topological data analysis and a measure theoretic formalism. The aim of this section is to introduce the mathematical background needed to define measure topological networks and the family of distances $d_{\mathrm{TpOT},p}$ described in the Introduction.
@@ 2.1 Persistent homology and persistent homology-hypergraphs
Let $(X,d)$ be a finite metric space – for example, a point cloud $X=\{x_{1},\cdots,x_{N}\}\subset\mathbb{R}^{n}$ as illustrated in Figure 2(a). For any $\varepsilon>0$, the Vietoris-Rips complex $K_{\varepsilon}(X)$ (see e.g. [21, Ch.III.2]) is the simplicial complex obtained from $X$ by adding a $k$-simplex $[x_{i_{0}},x_{i_{1}},\cdots,x_{i_{k}}]$ whenever the distance between all pairs of points in $\{x_{i_{0}},x_{i_{1}},\cdots,x_{i_{k}}\}$ is less than $\varepsilon$. Note that $K_{\varepsilon}(X)$ is a sub-complex of $K_{\varepsilon^{\prime}}(X)$ whenever $\varepsilon\leq\varepsilon^{\prime}$. Thus, as $\varepsilon$ grows, this yields a nested sequence of simplicial complexes:
$\mathcal{K}_{\rm VR}(X):=K_{\varepsilon_{0}}(X)\hookrightarrow K_{\varepsilon_{1}}(X)\hookrightarrow\cdots\hookrightarrow K_{\varepsilon_{M}}(X).$
@@
Here, the numbers $\varepsilon_{i}$, $i=1,\ldots,M$, are the parameters corresponding to the creation of new simplices; see Figure 2(b).
@@
Figure 2: Computing PH and the PH-hypergraph (a) A point cloud and (b) a filtration of simplicial complexes built on it. (c) The $1$-dimensional persistence diagram of the filtration in (b). (d) The PH-hypergraph constructed after computing a representative cycle (orange and purple cycles in (b)) for each class in (c).
@@
More generally, for a finite metric space $(X,d)$, a filtration over $X$ is a finite sequence $\mathcal{K}=(K_{\varepsilon_{1}},\ldots,K_{\varepsilon_{M}})$ of simplicial complexes, indexed by an increasing sequence of real numbers $\varepsilon_{i}\leq\varepsilon_{i+1}$, together with simplicial maps $K_{\varepsilon_{i}}\to K_{\varepsilon_{i+1}}$ for all $i$, such that the vertex set of each $K_{\varepsilon_{i}}$ is a subset of $X$. Common alternatives to the Vietoris-Rips complex in the literature with a similar interpretation include the Čech complex [21, Ch.III.2] and the Alpha complex [21, Ch.III.4]. Different constructions of the filtration $\mathcal{K}$ will influence the results of the persistent homology computation and their interpretation – our framework is general and does not depend on a specific filtration construction, as long as generator representatives can be retrieved computationally.
@@
A filtration $\mathcal{K}$ over $X$ can be analysed by computing the simplicial homology groups of each simplicial complex in the sequence and the corresponding induced maps. Collectively, this information is called the persistent homology (PH) of $\mathcal{K}$. The Structural Theorem [68] guarantees that, when using coefficients in $\mathbb{Z}/2\mathbb{Z}$, the birth, death, and evolution of homology classes in a filtration of simplicial complexes can be summarised in a multi-set called a persistence diagram. Let $\Lambda$ denote the upper-diagonal cone in the positive quadrant, $\Lambda=\{(x,y)\in\mathbb{R}^{2},y>x>0\}$. The persistence diagram $D$ associated to the filtration $\mathcal{K}$ is then a collection of points in $\Lambda$:
$D=\left\{(b_{i},d_{i})\right\}_{i=1}^{|D|}\>\subset\>\Lambda.$
@@
Each point $(b_{i},d_{i})\in D$ represents a homology class in the filtration $\mathcal{K}$, with coordinates corresponding to the birth $b_{i}$ and death $d_{i}$ parameters of the class (see Figure 2(c)). The persistence $d_{i}-b_{i}>0$ of a homology class is a measure of significance of the underlying topological feature. The persistence diagram $D$ can be interpreted as a topological descriptor of $X$. A local, geometric realisation of each homology class can be obtained by computing a representative cycle generating the class: see for instance the orange and purple cycles in Figure 2(b). Computing cycles depends on a number of choices [6, 18, 44, 60, 38], and therefore different representatives may lead to differing downstream results.
@@
The recently introduced PH-hypergraph [2] encodes the algebraic information of a persistence diagram, together with the local geometric interpretation of generating cycles, in a single higher-order network. Recall that a hypergraph $(V,E)$ consists of a vertex set $V$ and a set of hyperedges $E$, such that each hyperedge $e\in E$ is a subset of $V$, i.e. $e\subseteq V$. Given a point cloud $X$, we may construct an associated filtration $\mathcal{K}$ together with a choice $g=\{c_{1},\ldots,c_{|D|}\}$ of generating cycles for classes in the corresponding persistence diagram $D$. These data can be summarised in a PH-hypergraph $H=H(X,\mathcal{K},g)$ (see Figure 2(d)) in which we take $X$ to be the vertex set, and one hyperedge added for each cycle in $g$. Specifically, for every cycle $c\in g$, a point $x_{i}\in X$ is contained in the corresponding hyperedge $e(c)$ if it is a vertex of any simplex in $c$. While the hypergraph structure $H$ clearly depends on the choice of generators $g$ and the choice of filtration $\mathcal{K}$, for a fixed filtration $\mathcal{K}$ the qualitative (hyper)network structure of the PH-hypergraph has been empirically shown to be stable to noise and different choices of generating cycles [2].
@@ 2.2 Optimal transport on persistence diagrams
Given a Polish space $X$, let $\mathcal{P}(X)$ denote the set of probability measures supported on $X$. A coupling between two distributions $\mu,\mu^{\prime}\in\mathcal{P}(X)$ is a measure $\pi\in\mathcal{P}(X\times X)$ with left and right marginals equal to $\mu$ and $\mu^{\prime}$, respectively. That is,
$\pi(A\times X)=\mu(A)\quad\mbox{and}\quad\pi(X\times A)=\mu^{\prime}(A)$
for all measurable $A\subset X$. We denote the set of all couplings as $\Pi(\mu,\mu^{\prime})$.
Now let us equip $X$ with a metric $d$. For $p\in[1,\infty)$, the $p$-Wasserstein distance [61, Definition 6.1] between $\mu$ and $\mu^{\prime}$ is then defined as
$d_{\mathrm{W},p}(\mu,\mu^{\prime}):=\inf_{\pi\in\Pi(\mu,\mu^{\prime})}\left(\int_{X\times X}d(x,y)^{p}\>\mathrm{d}\pi(x,y)\right)^{1/p}.$ (1)
This metric is the central object of study in the optimal transportation literature [61, 51, 46]. It is a standard result that the infimum in (1) is realised (i.e. it is in fact a minimum) and a minimiser $\pi$ of (1) is referred to as an optimal coupling. Intuitively, this is a probabilistic matching between points in the supports of $\mu$ and $\mu^{\prime}$.
The optimal transport approach of finding matchings can be related to notions of distance on the space of persistence diagrams: the most popular construction is known as the bottleneck distance [14], which is related to the $\infty$-Wasserstein distance (although, we note that the initial introduction of the bottleneck distance made no reference to optimal transport). Our interest in this article is directed to similar versions of the $p$-Wasserstein distances for $1\leq p<\infty$. The usual modification of the Wasserstein matching problem to compare persistence diagrams is to allow points to be matched to a virtual point $\partial_{\Lambda}$, representing the diagonal set $\{(a,a)\mid a\in\mathbb{R}\}\subset\mathbb{R}^{2}$ (i.e. homology classes with vanishing persistence). Formally, for $p\in[1,\infty)$, the $p$-Wasserstein distance between two persistence diagrams $D$ and $D^{\prime}$ (see [37, Equation 6]) is defined as
$d_{\mathrm{W},p}^{\mathrm{PD}}(D,D^{\prime})=\min_{\pi\in\Pi(D,D^{\prime})}\left(\sum_{(a,b)\in\pi}\|a-b\|^{p}_{p}+\sum_{s\in U_{\pi}}\|s-\operatorname{Proj}_{\partial_{\Lambda}}(s)\|^{p}_{p}\right)^{1/p},$ (2)
@@
  • Slightly abusing notation, to illustrate the analogy to the definition of Wasserstein distance in (1), we denote by $\Pi(D,D^{\prime})$ the set of all partial matchings between points representing homology classes in $D$ and $D^{\prime}$. For any $\pi\in\Pi(D,D^{\prime})$, $\pi\subset D\times D^{\prime}$ has the property that each $a\in D$ and $b\in D^{\prime}$ appear in at most one ordered pair in $\pi$.
@@
The output from solving (2) is a partial matching $\pi$ between homology classes in $D$ and $D^{\prime}$ which is optimal with respect to the cost induced by the $\|\cdot\|_{p}$-norm in $\mathbb{R}^{2}$, restricted to $\Lambda$. By construction, this matching only depends on the coordinates of points in the persistence diagram. The resulting matching is thus agnostic to aspects of the data that are not captured in the persistence diagram representations. While the generality of the persistence diagram representation may contribute to its wide applicability, in many contexts this may be a limitation.
Observe that the notation and concepts used to define $d_{\mathrm{W},p}^{\mathrm{PD}}$ are similar to those that appear in the optimal transport framework described above; indeed, the family of persistence diagram distances (2) can be generalised via the language of partial OT [23, 19] by associating measures to the objects under study. We define a measure persistence diagram (MPD) to be a Radon measure $\mu$ supported on $\Lambda$. Then, any finite persistence diagram $D=\{(b_{i},d_{i})\}_{i=1}^{|D|}$ determines a measure persistence diagram: